OpenStudy (anonymous):

Solve each of these systems using the substitution method. Make sure to check your answers ! 2x + 3y = -4 x − 2y = 5 Answer x = 1, y = -2 x = -1, y = 2 x = 1, y = 2 x = -2, y = -1

OpenStudy (anonymous):

i cant see it.

OpenStudy (bibby):

uh x - 2y = 5

OpenStudy (anonymous):

x = 1, y= -2?

OpenStudy (bibby):

x=2y-5 2x + 3y = -4 2(2y-5) +3y = -4

OpenStudy (anonymous):

so the first one should be correct?

OpenStudy (triciaal):

@bibby 5 + 2y

OpenStudy (bibby):

oops

OpenStudy (bibby):

I should not be mathing right now

OpenStudy (triciaal):

@jjonty to use the substitution method using either of the equations find the expression for one of the variable in terms of the other then substitute in the other equation. for this one @bibby is using the expression for x from the second equation to substitute in the first. please note it should be 5 + 2y this gives one equation with one variable. after solving then substitute this value in either of the original equation to find the value of the other variable.

OpenStudy (anonymous):

oh ok thnx

OpenStudy (triciaal):

2(5 + 2y) + 3y = -4 10 + 4y + 3y = -4 y =

OpenStudy (anonymous):

-2

OpenStudy (triciaal):

good now you have y and you know x = 5 + 2y so x = 5 + 2*-2 = 5-4 =

OpenStudy (anonymous):

x = 1, y = -2

OpenStudy (triciaal):

you got it

OpenStudy (anonymous):

Ty

OpenStudy (triciaal):

make sure to check your answers you are welcome

OpenStudy (anonymous):

i will :)